Show that matrix A given below is skew-hermitian:


A=[0 -5-i -2+3i]

[ 5-i -3i -6]

[2+3i 6 6i ]
1
Expert's answer
2020-02-24T11:50:04-0500

The matrix "A" is skew-hermitian if it satisfies the relation


"A" skew-hermitian ⟺ "A^H = -A"


where "A^{H}"denotes the conjugate transpose of the matrix "A". In component form, this means that


"A" skew-hermitian ⟺ "a_{ij} = -\overline{a}_{ji}"


for all indices "i"  and "j" , where "a_{ij}" is the element in the "j"th row and "i"th column of "A" , and the overline denotes complex conjugation.


Therefore, to show that given matrix "A" is skew-hermitian, you should show that "A^H = -A".


"A = \begin{bmatrix}
 0 & -5-i & -2+3i\\
 5-i & -3i & -6 \\
 2+3i & 6 & 6i
\end{bmatrix}"


Determine "-A".


"-A = \begin{bmatrix}
 0 & 5+i & 2-3i\\
 -5+i & 3i & 6 \\
 -2-3i & -6 & -6i
\end{bmatrix}"


This equals complex conjugation:


"= \begin{bmatrix}
 \overline{0} &\overline{5-i} & \overline{2+3i}\\
 \overline{-5-i} & \overline{-3i} & \overline{6} \\
 \overline{-2+3i} & \overline{-6} & \overline{6i}
\end{bmatrix}"


It equals the transposed matrix:


"= \begin{bmatrix}
 \overline{0} &\overline{-5-i} & \overline{-2+3i}\\
 \overline{5-i} & \overline{-3i} & \overline{-6} \\
 \overline{2+3i} & \overline{6} & \overline{6i}
\end{bmatrix}^{T} = A^H"


Thus, the matrix "A" is skew-hermitian.


Use information from Wikipedia: https://en.wikipedia.org/wiki/Skew-Hermitian_matrix